1.If you deposit $100 in a bank account and the reserve ratio is 20 percent.

a.What is the minimum amount of money banks will be required to keep in reserves? How much loans can banks make at most? What is the money multiplier? How much money can be created from $100 of reserves?

b.If the fed raises the required reserve ratio to 30 percent. What is the minimum amount of money banks will be required to keep in reserves? How much loans can banks make at most? What is the money multiplier? How much money can be created from $100 of reserves?

2. How does the money supply change? Does it increase or decrease? What happens to the interest rate when the MS changes?
Answer #1

1. A) If deposit is $100 and bank's reserve ratio is 20% then bank will be required to keep $(100 * 20%) = $20 in reserves.

The bank can make a maximum of (100-20)% = 80% of the deposit $100 loan. So, the bank can make $80 loan at most.

Money multiplier = 1/reserve ratio = 1/20% = 5

The money created from this $100 deposit = $100 * money multiplier = $500.

B) If the FED raises the required reserve ratio to 30% then bank will be required to keep 30% of it's deposits as reserves. Therefore, of $100 deposit, the will be required to keep $(100*30%) = $30 as reserve.

The bank can make $(100 - 30) = $70 loan at most.

Money multiplier = 1/ required reserve ratio = 1/30% = 10/3.

Therefore, $[100* (10/3)] = $333.333 can be created from the $100 of reserves.